OpenStudy (anonymous):

A cylinder has a volume of 8000 cubic cm and base of 2000 cm. What is the height of the cylinder?

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

\[V = pir^2h\] plug in values for your volume and the base ( pir^2) then rearrange

OpenStudy (anonymous):

since we know that \[V=\pi r^2 h\] and \[\pi r^2 \]= surface area of a circle we take 8000/2000 to get 4cm

OpenStudy (anonymous):

I still dont understand, is this right?

OpenStudy (anonymous):

Two equations, two variables: \[ 8000 = \pi r^2h\\ 2000 = 2r \]

OpenStudy (mathstudent55):

By "base of 2000 cm" do you mean the base has an area of 2000 cm^2?

OpenStudy (anonymous):

Work backwards.

OpenStudy (anonymous):

V= Area * height , substitution you get 8000 = 2000* h 8000/2000 = h 4 = h